OpenStudy (kammykoo):

Juan analyzes the amount of radioactive material remaining in a medical waste container over time. He writes the function f(x) = 10(0.98)x to represent the amount of radioactive material that will remain after x hours in the container. Rounded to the nearest tenth, how much radioactive material will remain after 10 hours?

OpenStudy (kinged):

Substitute 10 for x.

OpenStudy (mathmale):

For the sake of clarity (which can be all important in problem solving), please use the symbol ^ to indicate exponentiation: Incorrect: f(x) = 10(0.98)x correct: f(x) = 10(0.98)^x, or

OpenStudy (mathmale):

\[f(x) = 10(0.98)^x\]

OpenStudy (mathmale):

Since the quantity inside parentheses, 0.98, is less than 1, the function f(x) is decreasing. I agree that you need only substitute 10 for x in this equation and then evaluate the result.